OpenStudy (anonymous):

Need in 10 minutes!! What are the classifications of each system? A. consistent independent B. coincident C. inconsistent 1. 4x + y = 8 x + 3y = 8 2. -4x + 6y = -2 2x - 3y = 1 3. 5x - 2y = 4 5x - 2y = 6 4. y = 3x - 4 -3x + y = -2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Babynini

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Can anyone help me please?

Nnesha (nnesha):

https://www.desmos.com/ use this

OpenStudy (astrophysics):

Well assuming you're not going to graph them..if you solve the system and get something like 0 = 8 that means there is no solution and is inconsistent. If you get 0 = 0 then there is a infinite amount of solutions and the system is dependent and a coincident. If you get something like x = 1, or something then there is one solution and the system is independent.

OpenStudy (astrophysics):

Or you can use desmos L

OpenStudy (anonymous):

I know those, thank you. I was just wondering if you could explain how to graph them?

OpenStudy (astrophysics):

You can set the equations up into y = mx+b form and graph them

Nnesha (nnesha):

 if the lines are parallel then there is no solution means (coincident) you can solve both equations for y you should get same slope different y-intercept

Nnesha (nnesha):

i meant to say inconsistent

OpenStudy (astrophysics):

Rather then us explaining it all just use this site, but if you still need help with graphing just ask http://www.algebra.com/algebra/homework/coordinate/Types-of-systems-inconsistent-dependent-independent.lesson

Nnesha (nnesha):

http://prntscr.com/94e3nw

OpenStudy (anonymous):

I just need help getting them into the equation y = mx + b.

OpenStudy (astrophysics):

Sure, so we have this equation 4x + y = 8, to set it up into y = mx+b form, notice we just solve for y. y=-4x+8

OpenStudy (astrophysics):

subtracted -4x on both sides, try the others it will be good practice

OpenStudy (anonymous):

So the second equation is x + 3y = 8. Would the slope-intercept form be: x + 3y = 8 3y = -x + 8 Then what would I do from there?

OpenStudy (anonymous):

@astrophysics

OpenStudy (astrophysics):

Exactly, but notice you have a 3y, so we need y on its own, so what can you do to get rid of that 3?

OpenStudy (anonymous):

Divide by 3.

OpenStudy (astrophysics):

Exactly, go ahead and do that

OpenStudy (anonymous):

How would you divide -x by 3?